OpenStudy (fearbigjjrob):

can anyone help me to factor out 4x^2+4x-3? I'm not sure how to get the end result, I know it is (2x-1)(2x+3) but I don't know how to work it out to get to that.

OpenStudy (fearbigjjrob):

ok I found out how to do it finally

OpenStudy (steve816):

Nice job then!

OpenStudy (mathmate):

@fearbigJJRob In case you're still interested in having a systematic way to factor the quadratic, you can do it as follows. The product of the first and last coefficient is 4*(-3)=-12 You look for two numbers whose product is -12 and sum +4 (middle coefficient). By trial and error with integers, you find +6 and -2. NOW, you rewrite the quadratic as: 4x^2\(\color{red}{+6}\)x \(\color{red}{-2}\)x-3 =2x(2x+3) - (2x+3) =(2x-1)(2x+3)